Differentiate the following functions:

(a) y = `4√x + 4/x + 3 ln x


1
Expert's answer
2021-05-07T14:26:22-0400

Let us differentiate the function

"y = 4\sqrt{x} +\frac{ 4}{x} + 3 \ln x:"

"y '= 4\frac{1}{2\sqrt{x}} -\frac{ 4}{x^2} + \frac{3 }{ x}= \frac{2}{\sqrt{x}} -\frac{ 4}{x^2} + \frac{3 }{ x}"